OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics. Center: (4, −4); solution point: (1, 0)

OpenStudy (anonymous):

(x-4)^2 + (y+4)^2 = r^2 now plug in x=1,y=0 and find r

OpenStudy (anonymous):

ok ?